# Constrained mean-field control with singular control: Existence, stochastic maximum principle and constrained FBSDE

### v1

@@ Abstract
This paper studies some mean-field control (MFC) problems with singular control under general dynamic state-control-law constraints. We first propose a customized relaxed control formulation to cope with the dynamic mixed constraints and establish the existence of an optimal control using some compactification arguments in the proper canonical spaces to accommodate the singular control. To characterize the optimal pair of regular and singular controls, we treat the controlled McKean-Vlasov process as an infinite-dimensional equality constraint and recast the MFC problem as an optimization problem on canonical spaces with constraints on Banach space, allowing us to derive the stochastic maximum principle (SMP) and a constrained BSDE using a novel Lagrange multipliers method. In addition, we further investigate the uniqueness and the stability result of the solution to the constrained FBSDE associated to the constrained MFC problem with singular control.
Keywords: Mean-field control with singular control, dynamic state-control-law constraints, relaxed control, compactification, stochastic maximum principle, Lagrange multipliers method, constrained FBSDE
@@ 1 Introduction
To overcome the high complexity in large stochastic system caused by the mutual influence among agents, mean-field game (MFG) problems (see [22] and [26]) has been proposed and popularized in academic studies, which is to find a Nash equilibrium in a competitive game formulation where each agent’s impact on the aggregate distribution is negligible. Recently, MFC problems has also gained considerable attention, which concerns the cooperative population where all agents share the same goal to attain the social optimum. The present paper targets at the MFC problem, which is often interpreted as an optimal control of McKean-Vlasov dynamics by a social planner. Some pioneer studies on this topic, just to name a few, can be found in [2], [7], [10], [1], [5] for the method of stochastic maximum principle (SMP); [25] and [15] for the limit theories; [27] and [30] for the dynamic programming principle (DPP) and the HJB equation on the Wasserstein space of probability measures; [4], [31] and [32] for the study of viscosity solution and comparison principle of the HJB equation. However, it is notable that most existing studies on MFC are limited to regular controls.
@@
Singular control, also called bounded variation stochastic control, allows the displacement of the states to be discontinuous. In single agent’s models, optimal singular control problems have been extensively investigated. For instance, we refer to [29] for early discussions on smooth fit principle; [20] for the existence of an optimal singular control based on the relaxed control formulation; [21] for the proof of DPP to account for singular controls; [3] for the stochastic maximum principle tailor-made for singular controls; [19] for some general theories on smooth fit principle. In the competitive game framework, there are a handful of studies on singular controls, see [11] for the n-player game with singular control as well as [17], [8] and [9] for some MFG problems with singular control. However, theoretical studies for MFC problems with singular control are underexplored in the literature, and only a recent paper [14] studies the derivation of DPP and the quasi-variational inequality using two-layer parameterisations of stochastic processes.
@@
The goal of the present paper is to contribute to the study of MFC problems by featuring both the singular control and the joint dynamic state-control-law constraints. Motivated by practical risk regulations, liquidity requirements, or feasibility restraints, different types of constraints have been considered in stochastic control problems, but predominantly in single agent’s model. Recently, in mean-field models with only regular controls, there are some emerging research to incorporate state and/or law constraints, see [18] for the level-set approach to tackle the law constraint in MFC problems; [12] for the analysis of mean-field limits under state constraints; [13] for the formulation of controlled Fokker-Planck equation under state constraints; [6] for the optimization formulation of the constrained MFC problem and the derivation of the stochastic maximum principle using the generalized Lagrange multipliers method.
@@
It is evident that the allowance of singular control in the mean-field model will entangle with the dynamic mixed constraints, which renders some existing methods for constrained MFC with regular controls not applicable in our framework. For example, it is difficult to generalize the level set approach in [18] to account for both control constraints and the singular control. Due to the possible discontinuity of the state process and its law under the singular control, characterizing the associated Fokker-Planck equation is still an open problem in the literature, and hence the approach in [13] is also not suitable to handle the singular control. On the other hand, some conventional methods that have been employed in unconstrained MFG and MFC problems with singular control such as DPP and compactification arguments also need to be carefully and substantially modified in the setting when pathwise constraints are imposed jointly on the state process and its distribution flow.
@@
The present paper fills the gap in the literature by advancing some theoretical foundations for constrained MFC problems with singular control. The first contribution of the paper is to address the existence of an optimal pair of regular and singular controls by developing the customized relaxed control formulation and martingale problem to embed the dynamic mixed constraints. That is, we embed the dynamic mixed constraints in item $(iv)$ of the admissible control rules in Definition 3.1. We then present in Lemma 3.5 a martingale problem characterization associated to the admissible control rules and establish in Lemma 3.7 the closedness property of the set of probability measures fulfilling the constraints. By developing some technical compactification arguments to accommodate the singular control and combining the previous lemmas using the embedded constraints, we are able to show that the set of approximately optimal control rules is compact (see Lemma 3.8) and hence the set of optimal control rules is compact and convex that yields the existence of an optimal relaxed control (see Theorem 3.9). By some standard arguments, we also obtain the existence of an optimal strict control for the constrained MFC problem with singular control.
@@
As the third contribution of the present paper, we also analyze the non-standard constrained FBSDE in (26)-(27) associated to our constrained MFC problem with singular control. It is worth noting that the Lagrange multiplier $\eta$ in Theorem 4.4 is treated as part of the solution to the constrained BSDE problem. We then verify the constraints qualification and obtain the KKT optimality condition in Lemma 5.1. As a result, we further establish the uniqueness of the solution to the nonstandard constrained FBSDE in (26)-(27), see Theorem 5.3. Furthermore, under some enhanced assumptions, we also prove the stability result of the solution in the sense of (31) in Theorem 5.5.
The rest of the paper is organized as follows. Section 2 introduces the constrained MFC problem with singular control and some standing assumptions imposed on the underlying model. In Section 3, we introduce the relaxed control formulation and establish the existence of optimal control to the constrained MFC problem by applying some compactification arguments. Section 4 formulates the equivalent optimization problem with constraints in Banach space, for which we develop the SMP and ajoint constrained BSDE using Lagrange multipliers method. Finally, the uniqueness and the stability result of the solution to the constrained FBSDEs associated to the constrained MFC problem with singular control are discussed in Section 5.
Notations. For the reader’s convenience, we list below some notations that will be frequently used throughout the paper:
$(E^{*},\lVert\cdot\rVert_{E^{*}})$ The dual space of a Banach space $(E,\lVert\cdot\rVert_{E})$
$\langle\cdot,\cdot\rangle_{E,E^{*}}$ The pairing between Banach space $(E,\lVert\cdot\rVert_{E})$ and its dual
$a\geq_{K}0$ ($a>_{K}0$) $a\in K$ ($a\in K^{\rm o}$) with $K$ being a closed (linear) cone
$Df$ The Fréchet derivative of $f:E\mapsto\mathbb{R}$
$\nabla_{x}f$ Gradient of $f:\mathbb{R}^{n}\mapsto\mathbb{R}$
$D_{X}f$ The partial Fréchet derivative of $f$ with respect to $X$
$D_{\alpha}f$ The partial Fréchet derivative of $f$ with respect to $\alpha$
$L^{p}((A,\mathscr{B}(A),\lambda_{A});E)$ Set of $L^{p}$-integrable $E$-valued mapping defined on
measure space $(A,\mathscr{B}(A),\lambda_{A})$. Write $L^{p}(A;E)$ for short
$L^{p}((A,\mathscr{B}(A),\lambda_{A});E)/\mathord{\sim}$ The quotient set of $L^{p}(A;E)$ w.r.t. equivalence
[…]
@@ 2.1 MFC with singular control and mixed constraints
For the above controlled McKean-Vlasov dynamics, we shall impose some dynamic state-control-law constraints. To this end, let $d\in\mathbb{N}$ be the number of inequality constraints. For $i\in I:=\{1,\dots,d\}$, let us consider $\phi^{i}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\mapsto\mathbb{R}$ with $i\in I$ such that the dynamic state-control-law inequality constraints are described by
$\phi^{i}(t,X_{t}^{\alpha,\zeta},\mathcal{L}(X_{t}^{\alpha,\zeta}),\alpha_{t})\geq 0,~{}\mathrm{d}t\times\mathrm{d}\mathbb{P}\text{-}\mathrm{a.s.},$ (2)
where $X^{\alpha,\zeta}=(X_{t}^{\alpha,\zeta})_{t\in[0,T]}$ obeys the McKean-Vlasov SDE (1). We denote $\mathscr{U}_{\rm ad}[0,T]$ as the admissible control set of processes $(\alpha,\zeta)\in\mathscr{U}[0,T]\times\mathcal{A}^{k}[0,T]$ such that there exists a cádlág $\mathbb{F}$-adapted process $\mathbb{R}^{n}$-valued $X^{\alpha,\zeta}=(X_{t}^{\alpha,\zeta})_{t\in[0,T]}$ satisfying (1) and (2).
@@
where $f:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\mapsto\mathbb{R}$ and $c:[0,T]\mapsto\mathbb{R}^{k}$ are respectively the running cost function and the cost function associated with singualr controls.
@@ 2.2 Standing assumptions
    The mappings $b,\sigma,f$ and $\phi^{i}$ for $i\in I$ are Borel measurable, $b,\sigma$ are continuous and $f$ is lower semi-continuous (l.s.c.) in $(x,\mu,u)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$.
@@
    For $i\in I$, $\phi^{i}$ is upper semi-continuous (u.s.c.) in $(x,\mu,u)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$ and there exists a constant $M>0$ such that, for all $i\in I$ and $(x,u)\in\mathbb{R}^{n}\times\mathbb{R}^{l}$,
    $|\phi^{i}(t,x,\mu,u)|\leq M(1+|x|+M_{2}(\mu)+|u|)$
    with $M_{2}(\mu):=\int_{\mathbb{R}^{n}}|x|^{2}\mu(\mathrm{d}x)$. Moreover, $\phi^{i}$ is uniformly u.s.c. in $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$ with respect to $(t,x,u)\in[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}$.
@@
    The mappings $b(t,x,\mu,u)$ and $\sigma(t,x,\mu,u)$ are uniformly Lipschitz continuous in $(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$ in the sense that, there exists a constant $M>0$ independent of $u\in\mathbb{R}^{l}$ such that, for all $(x,\mu),(x^{\prime},\mu^{\prime})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$ and $(t,u)\in[0,T]\times\mathbb{R}^{l}$,
    $\left|b(t,x^{\prime},\mu^{\prime},u)-b(t,x,\mu,u)\right|+\left|\sigma(t,x^{\prime},\mu^{\prime},u)-\sigma(t,x,\mu,u)\right|\leq M(|x-x^{\prime}|+\mathcal{W}_{2,\mathbb{R}^{n}}(\mu,\mu^{\prime})),$
    where $\mathcal{W}_{2,\mathbb{R}^{n}}$ is a $2$-Wasserstein metric on $\mathcal{P}_{2}(\mathbb{R}^{n})$.
@@
    There exists a constant $M>0$ such that, for all $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$,
    $|b(t,x,\mu,u)|+|\sigma(t,x,\mu,u)|\leq M(1+|x|+M_{2}(\mu)+|u|).$
@@
    $G:[0,T]\mapsto\mathbb{R}^{n\times k}$ and $c:[0,T]\mapsto[0,\infty)^{k}$ are deterministic continuous functions. Moreover, $c(t)=(c^{1}(t),\ldots,c^{k}(t))$ satisfies $c^{i}(t)\geq c_{0}>0$ for all $t\in[0,T]$.
@@
    For $i\in I$, the mappings $b,\sigma,f,g$ and $\phi^{i}$ are Borel measurable and continuously differentiable w.r.t. $x\in\mathbb{R}^{n}$. Moreover, $\phi^{i}$ is jointly continuous in $(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$.
@@
    There exists a constant $M>0$ such that, for all $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$,
@@
    There exists a constant $M>0$ such that, for all $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$, $|\delta(t,x,\mu,u)|\leq M(1+|x|+|u|+M_{2}(\mu))$, where $\delta\in\{b,\sigma,\nabla_{u}b(\cdot),\nabla_{u}\sigma(\cdot),\nabla_{x}f(\cdot),\nabla_{u}f(\cdot)\}$.
@@
    For $i\in I$, the mappings $b(t,x,\mu,u),\sigma(t,x,\mu,u),f(t,x,\mu,u)$ and $\phi^{i}(t,x,\mu,u)$ are continuously L-differentiable w.r.t. $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$. Moreover, there exists a constant $M>0$ such that, for all $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$,
    $\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\mu}f(t,x,\mu,u)(x^{\prime},u^{\prime})|^{2}\mu(\mathrm{d}x^{\prime},\mathrm{d}u^{\prime})\right)^{\frac{1}{2}}\leq M(1+|x|+|u|+M_{2}(\mu)),$
    and moreover
    $\sum_{i\in I}\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\mu}\phi^{i}(t,x,\mu,u)(x^{\prime},u^{\prime})|^{2}\mu(\mathrm{d}x^{\prime},\mathrm{d}u^{\prime})\right)^{\frac{1}{2}}\leq M.$
@@
    The constrained function $\phi$ is independent of the control variable and has the linear form of
    $\phi(t,x,\mu)=A(t)x+B(t)\int_{\mathbb{R}^{n}}x\mu(\mathrm{d}x)+C(t),\quad\forall(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}),$
    where $A=(A(t))_{t\in[0,T]}$, $B=(B(t))_{t\in[0,T]}$ and $C=(C(t))_{t\in[0,T]}\in C([0,T];\mathbb{R}^{n})$ and it holds that $|A(t)|>0$ and $|A(t)+B(t)|>0$ for all $t\in[0,T]$.
@@
    The Hamiltonian $H$ defined by $H(t,x,\mu,u,y,z):=b(t,x,\mu,u)^{\top}y+\mathrm{tr}(\sigma(t,x,\mu,u)^{\top}z)+f(t,x,\mu,u)$ is strictly convex in $u\in U$. There exists a measurable mapping $\hat{u}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto U\subset\mathbb{R}^{l}$ with at most quadratic growth in $(x,\mu,y,z)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ such that $\hat{u}(t,x,\mu,y,z)\in\mathop{\arg\min}_{u\in U}H(t,x,\mu,u,y,z)$ and $\hat{u}$ is Lipschitz continuous in the sense that there exists a constant $M>0$ such that, for all $(x_{i},\mu_{i},y_{i},z_{i})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ with $i=1,2$,
@@
    We then set $\hat{b}(t,x,\mu,y,z):=b(t,x,\mu,\hat{u}(t,x,\mu,y,z))$, $\hat{\sigma}(t,x,\mu,y,z):=\sigma(t,x,\mu,\hat{u}(t,x,\mu,y,z))$ and $\hat{f}(t,x,\mu,y,z):=f(t,x,\mu,\hat{u}(t,x,\mu,y,z))$.

### v2

@@ Abstract
This paper studies a class of mean-field control (MFC) problems with singular control under general dynamic state-control-law constraints. We first propose a customized relaxed control formulation to cope with the dynamic mixed constraints and establish the existence of an optimal control using compactification argument in the proper canonical spaces to accommodate the singular control. To characterize the optimal pair of regular and singular controls, we treat the controlled McKean-Vlasov process as an infinite-dimensional equality constraint and recast the MFC problem as an optimization problem on canonical spaces with constraints on Banach space, allowing us to derive the stochastic maximum principle (SMP) and a constrained BSDE using a novel Lagrange multipliers method. Additionally, we investigate the uniqueness and the stability result of the solution to the constrained FBSDE associated to the constrained MFC problem with singular control.
Keywords: Mean-field control, singular control, dynamic state-control-law constraint, compactification, stochastic maximum principle, Lagrange multiplier, constrained FBSDE
@@ 1 Introduction
To overcome the high complexity in large stochastic system caused by the mutual influence among agents, mean-field game (MFG) problems (see [22] and [26]) has been proposed and popularized in academic studies, which is to find a Nash equilibrium in a competitive game formulation where each agent’s impact on the aggregate distribution is negligible. MFC problems has also gained considerable attention, which concerns the cooperative population where all agents share the same goal to attain the social optimum. The present paper targets at the MFC problem, which is often interpreted as an optimal control of McKean-Vlasov dynamics by a social planner. Some pioneer studies on this topic, just to name a few, can be found in [2, 7, 10, 1, 5] for the SMP method; [25] and [15] for the limit theory; [27] and [30] for the DPP and HJB equation on the Wasserstein space of probability measures; [4], [31] and [32] for the viscosity solution and comparison principle of the HJB equation. However, it is notable that most existing studies on MFC are limited to regular controls.
@@
Singular control, also called bounded variation stochastic control, allows the displacement of the states to be discontinuous. In single agent’s models, optimal singular control problems have been extensively investigated. For instance, we refer to [29] for discussions on smooth fit principle; [20] for the existence of optimal singular controls based on the relaxed control formulation; [21] for the proof of DPP to account for singular controls; [3] for the SMP tailor-made for singular controls; [19] for a general theory on smooth fit principle. In the competitive game framework, there are a handful of studies on singular controls, see [11] for the n-player game with singular control as well as [17, 8, 9] for some MFG problems with singular control. However, theoretical studies for MFC problems with singular control are underexplored in the literature, and only a recent paper [14] studies the derivation of DPP and the quasi-variational inequality using two-layer parameterisations of stochastic processes.
@@
The goal of the paper is to contribute to the study of MFC problems by featuring both the singular control and the joint dynamic state-control-law constraints. Motivated by practical risk regulations, liquidity requirements, or feasibility restraints, different types of constraints have been considered in stochastic control problems, but predominantly in single agent’s model. Recently, in mean-field models with only regular controls, there are some emerging research to incorporate state and/or law constraints, see [18] for the level-set approach to tackle the law constraint in MFC problems; [12] for the analysis of mean-field limits under state constraints; [13] for the formulation of controlled FP equation under state constraints; [6] for the optimization formulation of the constrained MFC problem and the derivation of the SMP using the generalized Lagrange multipliers method.
@@
It is evident that the allowance of singular control in the mean-field model will entangle with the dynamic mixed constraints, which renders some existing methods for constrained MFC with regular controls not applicable in our framework. For example, it is difficult to generalize the level set approach in [18] to account for both control constraints and the singular control. Due to the possible discontinuity of the state process and its law under the singular control, characterizing the associated FP equation is still an open problem in the literature, and hence the approach in [13] is also not suitable to handle the singular control. On the other hand, some conventional methods that have been employed in unconstrained MFG and MFC problems with singular control such as DPP and compactification arguments also need to be carefully and substantially modified in the setting when pathwise constraints are imposed jointly on the state process and its distribution flow.
@@
The present paper fills the gap in the literature by advancing some theoretical foundations for constrained MFC problems with singular control. The first contribution of the paper is to address the existence of an optimal pair of regular and singular controls by developing the customized relaxed control formulation and martingale problem to embed the dynamic mixed constraints. That is, we embed the dynamic mixed constraints in item (iv) of the admissible control rules in Definition 3.1. We then present in Lemma 3.5 a martingale problem characterization associated to the admissible control rules and establish in Lemma 3.7 the closedness property of the set of probability measures fulfilling the constraints. By developing technical compactification arguments to accommodate the singular control and combining the previous lemmas using the embedded constraints, we are able to show that the set of approximately optimal control rules is compact (see Lemma 3.8), and hence the set of optimal control rules is compact and convex that yields the existence of an optimal relaxed control (see Theorem 3.9). By standard argument, we also obtain the existence of an optimal strict control for the constrained MFC problem with singular control.
@@
As the third contribution of the present paper, we also analyze the non-standard constrained FBSDE in (5)-(27) associated to our constrained MFC problem with singular control. It is worth noting that the Lagrange multiplier $\eta$ in Theorem 4.2 is treated as part of the solution to the constrained BSDE problem. We then verify the constraints qualification and obtain the KKT optimality condition in Lemma 5.1. As a result, we further establish the uniqueness of the solution to the nonstandard constrained FBSDE in (5)-(27), see Theorem 5.4. Furthermore, under some enhanced assumptions, we also prove the stability result of the solution in the sense of (5.6) in Theorem 5.6.
The rest of the paper is organized as follows. Section 2 introduces the constrained MFC problem with singular control and standing assumptions imposed on the model. Section 3 gives the relaxed control formulation and establishes the existence of optimal control to the constrained MFC problem by applying compactification arguments. Section 4 formulates the equivalent optimization problem with constraints in Banach space, for which we develop the SMP and ajoint constrained BSDE using Lagrange multipliers method. Lastly, the uniqueness and the stability result of the solution to the constrained FBSDE associated to the constrained MFC problem with singular control are discussed in Section 5. The proofs of some auxiliary results are provided in Appendix.
Notations. We list below some notations that will be frequently used throughout the paper:
$(E^{*},\lVert\cdot\rVert_{E^{*}})$ Dual space of a Banach space $(E,\lVert\cdot\rVert_{E})$
$\langle\cdot,\cdot\rangle_{E,E^{*}}$ Pairing between Banach space $(E,\lVert\cdot\rVert_{E})$ and its dual
$a\geq_{K}0~(a>_{K}0)$ $a\in K~(a\in K^{\mathrm{o}})$, with $K$ a closed (linear) cone
DfDf Fréchet derivative of $f:E\to\mathbb{R}$
$\nabla_{x}f$ Gradient of $f:\mathbb{R}^{n}\to\mathbb{R}$
$D_{X}f$ Partial Fréchet derivative of $f$ with respect to $X$
$D_{\alpha}f$ Partial Fréchet derivative of $f$ with respect to $\alpha$
$L^{p}((A,\mathscr{B}(A),\lambda_{A});E)$ Set of $L^{p}$-integrable $E$-valued mappings on measure space
$(A,\mathscr{B}(A),\lambda_{A})$; write $L^{p}(A;E)$ for short
$L^{p}((A,\mathscr{B}(A),\lambda_{A});E)/\mathord{\sim}$ Quotient set of $L^{p}(A;E)$ under relation $f\mathord{\sim}g$ iff $f=g$ $\lambda_{A}$-a.e.;
@@ 2.1 MFC with singular control and mixed constraints
For the above controlled McKean-Vlasov dynamics, we introduce dynamic state-control-law constraints. Let $d\in\mathbb{N}$ be the number of inequality constraints. For $i\in I:=\{1,\dots,d\}$, consider $\phi^{i}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\mapsto\mathbb{R}$ with $i\in I$ such that the dynamic state-control-law inequality constraints are described by, $\mathrm{d}t\otimes\mathrm{d}\mathbb{P}$-a.s.
$\phi^{i}(t,X_{t}^{\alpha,\zeta},\mathcal{L}(X_{t}^{\alpha,\zeta}),\alpha_{t})\geq 0,$ (2)
where $X^{\alpha,\zeta}=(X_{t}^{\alpha,\zeta})_{t\in[0,T]}$ obeys the McKean-Vlasov SDE (1). Denote by $\mathscr{U}_{\rm ad}[0,T]$ the admissible control set of processes $(\alpha,\zeta)\in\mathscr{U}[0,T]\times\mathcal{A}^{k}[0,T]$ such that there exists a càdlàg $\mathbb{F}$-adapted ($\mathbb{R}^{n}$-valued) process $X^{\alpha,\zeta}=(X_{t}^{\alpha,\zeta})_{t\in[0,T]}$ satisfying (1) and (2).
@@
where $f:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\mapsto\mathbb{R}$ and $c:[0,T]\mapsto\mathbb{R}^{k}$ are respectively the running cost function and the cost function associated with singular controls.
@@ 2.2 Standing assumptions
    The mappings $b,\sigma,f$ and $\phi^{i}$, for $i\in I$ are Borel measurable, $b,\sigma$ are continuous and $f$ is lower semi-continuous (l.s.c.) in $(x,\mu,u)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$.
@@
    For $i\in I$, $\phi^{i}$ is upper semi-continuous (u.s.c.) in $(x,\mu,u)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$, and there exists $M>0$ such that $|\phi^{i}(t,x,\mu,u)|\leq M(1+|x|+\sqrt{M_{2}(\mu)}+|u|)$ for $(i,x,u)\in I\times\mathbb{R}^{n}\times\mathbb{R}^{l}$ and $M_{2}(\mu):=\int_{\mathbb{R}^{n}}|x|^{2}\mu(\mathrm{d}x)$. Moreover, $\phi^{i}$ is uniformly u.s.c. in $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$ w.r.t. $(t,x,u)\in[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}$.
@@
    The mappings $b(t,x,\mu,u)$ and $\sigma(t,x,\mu,u)$ are uniformly Lipschitz continuous in $(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$, i.e., there is $M>0$ independent of $u\in\mathbb{R}^{l}$ such that, for all $(x,\mu),(x^{\prime},\mu^{\prime})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$ and $(t,u)\in[0,T]\times\mathbb{R}^{l}$,
    $\left|b(t,x^{\prime},\mu^{\prime},u)-b(t,x,\mu,u)\right|+\left|\sigma(t,x^{\prime},\mu^{\prime},u)-\sigma(t,x,\mu,u)\right|\leq M(|x-x^{\prime}|+\mathcal{W}_{2,\mathbb{R}^{n}}(\mu,\mu^{\prime})).$
@@
    There is $M>0$ such that $|b(t,x,\mu,u)|+|\sigma(t,x,\mu,u)|\leq M(1+|x|+\sqrt{M_{2}(\mu)}+|u|)$ for $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$.
@@
    $G:[0,T]\mapsto\mathbb{R}^{n\times k}$ and $c:[0,T]\mapsto[0,\infty)^{k}$ are deterministic continuous functions. Moreover, there is $c_{0}>0$ such that $c^{i}(t)\geq c_{0}$ for all $t\in[0,T]$.
@@
    The mappings $b,\sigma,f$ and $\phi^{i}$ are Borel measurable and continuously differentiable w.r.t. $x\in\mathbb{R}^{n}$. Moreover, $\phi^{i}$ is jointly continuous in $(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$.
@@
    There exists $M>0$ such that, for all $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$,
@@
    There exists $M>0$ such that $|\delta(t,x,\mu,u)|\leq M(1+|x|+|u|+\sqrt{M_{2}(\mu)})$ for $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$ and $\delta\in\{b,\sigma,\nabla_{u}b(\cdot),\nabla_{u}\sigma(\cdot),\nabla_{x}f(\cdot),\nabla_{u}f(\cdot)\}$.
@@
    The mappings $b(t,x,\mu,u),\sigma(t,x,\mu,u),f(t,x,\mu,u)$ and $\phi^{i}(t,x,\mu,u)$ are continuously Lions-differentiable w.r.t. $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$. Moreover, there exists $M>0$ such that, for all $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$,
    $\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\mu}f(t,x,\mu,u)(x^{\prime})|^{2}\mu(\mathrm{d}x^{\prime})\right)^{\frac{1}{2}}\leq M\left(1+|x|+|u|+\sqrt{M_{2}(\mu)}\right),$
    and moreover $\sum_{i\in I}\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\mu}\phi^{i}(t,x,\mu,u)(x^{\prime})|^{2}\mu(\mathrm{d}x^{\prime})\right)^{\frac{1}{2}}\leq M$.
@@
    The constrained function $\phi$ is independent of the control variable and has the linear form $\phi(t,x,\mu)=A(t)x+B(t)\int_{\mathbb{R}^{n}}x\mu(\mathrm{d}x)+C(t)$ for $(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$. Here, $A=(A(t))_{t\in[0,T]}$, $B=(B(t))_{t\in[0,T]}$ and $C=(C(t))_{t\in[0,T]}\in C([0,T];\mathbb{R}^{n})$ and $|A(t)|>0$ and $|A(t)+B(t)|>0$ for $t\in[0,T]$.
@@
    The Hamiltonian $H$ defined by $H(t,x,\mu,u,y,z):=b(t,x,\mu,u)^{\top}y+\mathrm{tr}(\sigma(t,x,\mu,u)^{\top}z)+f(t,x,\mu,u)$ is strictly convex in $u\in U$. There exists a measurable mapping $\hat{u}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto U\subset\mathbb{R}^{l}$ with at most quadratic growth in $(x,\mu,y,z)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ such that $\hat{u}(t,x,\mu,y,z)\in\mathop{\arg\min}_{u\in U}H(t,x,\mu,u,y,z)$ and $\hat{u}$ is Lipschitz continuous, i.e., there is $M>0$ such that, for $(x_{i},\mu_{i},y_{i},z_{i})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ with $i=1,2$,
@@
    $\qquad\qquad\leq M(|x_{1}-x_{2}|+\mathcal{W}_{2,\mathbb{R}^{n}}(\mu_{1},\mu_{2})+|y_{1}-y_{2}|+|z_{1}-z_{2}|).$
    Set $\hat{b}(t,x,\mu,y,z):=b(t,x,\mu,\hat{u}(t,x,\mu,y,z))$, $\hat{\sigma}(t,x,\mu,y,z):=\sigma(t,x,\mu,\hat{u}(t,x,\mu,y,z))$ and $\hat{f}(t,x,\mu,y,z):=f(t,x,\mu,\hat{u}(t,x,\mu,y,z))$.